# FOOL: Addressing the Downlink Bottleneck in Satellite Computing with Neural Feature Compression

### v1

@@ 3.3.1 The Effectiveness of Shallow Features
Figure 3 visualizes the trade-off. We extract the features from a ResNet network with weights trained on the ImageNet [37] classification task and recover the original image by training separate reconstruction models for each marked location. Section 5.6 elaborates on the reconstruction, but here it is sufficient to view it as a black box function. Notice that models at shallow layers can recover the input with high similarity, but the recovery progressively worsens as the path distance increases. Now assume a discriminative model (e.g., a ResNet for image classification) $\mathcal{M}=(\mathcal{H},\mathcal{T})$ with separate shallow $\mathcal{H}$ from deeper layers $\mathcal{T}$ as disjoint subsets, such that $\mathcal{H}(X)=H$ (i.e, mapping input to shallow features) and $\mathcal{M}(X)=\mathcal{T}(\mathcal{H}(X))$. Further, given a codec $\mathrm{c}=(\mathrm{enc},\mathrm{dec})$ where $\mathrm{dec}(\mathrm{enc}(X))=\tilde{H}$ is an approximation of $\mathcal{H}(X)$. Then, $\tilde{H}$ is a sufficient approximation of $H$ if $\mathcal{T}(\tilde{H})$ results in lossless prediction, i.e., no drop in prediction performance relative to $\mathcal{T}(H)$. In other words, a sufficient representation in the shallow latent space results in high similarity in the deep latent space between $\mathcal{T}(H)$ and $\mathcal{T}(\tilde{H})$.
Consider that similarity in the deep latent space coincides with high similarity for human perception in the input space [18]. Therefore, the encoder output $\mathrm{enc}(X)$ (i.e., a compressed representation of the shallow features) should retain sufficient information to reconstruct $X$ with quality comparable to $\mathcal{H}(X)$ (exemplified in Figure 3). Finally, since $enc(X)$ sufficiently approximates $H$, it should be possible to sufficiently approximate the shallow representation of any model $\mathcal{M}^{\prime}=(\mathcal{H}^{\prime},\mathcal{T}^{\prime})$ if $I(\mathcal{H}(X),X)\approx I(\mathcal{H}^{\prime}(X),X)$.
3.3.2 Task Specificity over Target Perception for Rate Gains
We define perception as the input interface, i.e., the arrangement of available information an entity requires to consume a message. Human perception requires images in common formats and color spaces. Conversely, machine perception requires image or feature tensors. A common justification for task-oriented communication is that there are potential rate gains in machine-to-machine communication [47]. The idea is that messages for machine perception may require less information than human perception, i.e., that it should be possible to reduce bitrate by not (exclusively) using input reconstruction as the distortion measure. We argue that the framing is not entirely correct for the case without prior information on any downstream tasks.
Specifically, there is a contradiction in models outperforming human experts in various image-related tasks, i.e., machines can detect signals and patterns that humans (physiologically or intellectually) cannot. Hence, the opposite may hold, i.e., when compressing for human perception, we should see more potential for rate savings, not less. The claim is consistent with the results in Figure 1 where codecs with high reconstruction quality result in images that are deceptively similar to the input (details in Section 6.3). Consider the remaining predictive loss after adjusting weights to infer recoverable signals and input perturbations from applying a lossy image codec that reconstructs images that are difficult to distinguish from the original with human perception. Then, it is reasonable to assume that the rate gains relative to the uncompressed size are from signals crucial for models to outperform human experts. If the objective is a complete generalization for any possible downstream task on a dataset, lossless encoding will be the only viable option under most circumstances. However, once we relax the requirements as minimizing classification error for a constrained set of tasks, the compression model may achieve lossless prediction and trade-off distortion (e.g., PSNR) with lower rates. Conclusively, rate gains are from task specifity irrespective of the input interface, whether it is a particular layer of a DNN architecture, human receptors, or textual encoding. Note that this is irrespective of whether we constrain the solution space to discriminative tasks without involving any image reconstruction objective.
Besides visualizing Equation 4, the input image illustrates a practical example. The frog subset of ImageNet distinguishes between Tree Frogs, Bullfrogs, and Tailed Frogs. Since these frog species have distinct figures and primary colors, the more delicate characteristics of a tree frog are redundant for ImageNet classification. The network gradually discards information regarding the fine-grained blue-yellow colored patterns, permitting only the recovery of general shape and environment from the deep features. The deeper the features, the less structure and detail are present, which may be redundant for the task. Contrastingly, reconstruction from shallow layers results in images with high perceptual similarity to the input.
Now, suppose training a codec where the encoder retains the minimal information necessary to reconstruct the output of the deepest layers for a classification task (e.g., similar to Vista [24]). Then, we can reduce the transfer cost to as low as $\log_{2}{({\#\text{labels}})}$ without predictive loss. However, we may lack the information for other tasks, i.e., there is a trade-off between generalization and the lower bound on the bitrate. In contrast, targeting shallow features for compression may strike a balance between aiming to retain information for all possible downstream tasks and only emphasizing the salient regions for the tasks associated with a foundational model. Arguably, the limitation is negligible as maintainers will train foundational models to provide clients with valuable services.
@@ 4.1 Compression and Prediction Request Flow
Figure 4 provides a high-level overview for serving client requests.
@@
For samples processed by FOOL, there is a single encoder and its output $\boldsymbol{\hat{{\mathrm{y}}}}$ is forwarded to the detection pipeline, skipping the shallow layers. A decoder transforms $\boldsymbol{\hat{{\mathrm{y}}}}$ into a target representation. A predictor outputs bounding boxes for a specified task. An image reconstruction model optionally restores the latent to a human-interpretable image to overlay the bounding boxes. Samples downlinked with bent pipe or some image codec are forwarded to the shallow layers, skipping the FOOL decoder and reconstruction model. This section focuses on the pipeline, before Section 5 introduces the compression method.
@@ 4.2 Profiling Compression Pipelines for OEC
A common challenge for operators is to determine whether reported performance regarding resource usage or throughput from the latest advancements generalizes to their target hardware. This problem stems not from a lack of rigor by authors but from the sheer heterogeneity of the AI accelerator landscape [48]. Graph compilers and other vendor-specific optimizations (e.g., TensorRT [^2] https://developer.nvidia.com/tensorrt, Apache TVM [^3] https://tvm.apache.org) further complicate evaluation, with varying methods for operator fusion, graph rewriting, etc. Consequently, FOOL includes a simple profiling and evaluation strategy that operators may run before deployment (i.e., on the ground). Notably, in contrast to existing work that partitions images to match the input size of a particular application, the profiler regards the importance of spatial dimensions for resource efficiency. The purpose of the profiler is to determine a configuration that maximizes throughput. While throughput evaluation is straightforward, how to measure it (e.g., images/second) is not necessarily obvious, particularly for (neural) compression pipelines. Specifically, we did not find prior work to consider two subtle but significant differences between terrestrial and LEO remote sensing for throughput calculation.
First, terrestrial and LEO remote sensing with constrained sensor networks demand resource-conscious methods, but in LEO, downlinks are only available within segments. Due to memory and storage constraints, devices must process samples according to a sensor rate, i.e., a prolonged (around 1-30 seconds) interval between incoming samples. Hence, the objective in LEO is to maximize the number of pixels the accelerator can process before reaching a downlink segment, given a time constraint for a single sample (i.e., “frame deadline” [27]). For example, assume a cheaper and a costlier compression model where both models meet the frame deadline. Applying the latter results in half the bitrate but thrice the inference time. Using the former is beneficial in most network conditions for real-time terrestrial applications since it results in a lower end-to-end request latency. In contrast, applying the latter in LEO may be advantageous, as finishing earlier results in the needless idle time of resources.
Second, satellite imagery has substantially higher resolution than captures from most terrestrial sensor networks. A standard method to improve throughput for high-dimensional images is parallel processing with tile partitioning. The distinction is that there is more control over the spatial dimensions and the batch size. Nonetheless, a caveat is the friction between a model’s size and the input size. Increasing the width (e.g., the number of feature maps output by a convolutional layer) of a neural codec’s parametric transforms may result in better compression performance but lower processing throughput. In summary, we require a measure that includes (i) the tile spatial dimensions, (ii) batch size, and (iii) the capacity-compression performance trade-off.
@@
Notice how evaluating img/s always favors smaller spatial dimensions and disregards batch size and model width. In contrast, PP/s reveals that the optimal spatial dimension is around $500\times 500$ but will naturally favor smaller models, as it does not consider that wider models may reduce transfer costs. To alleviate the limitations of PP/s, we propose measuring Transfer Cost Reduction per Second (TCR/s) as:
@@
The measure now includes the compression performance as the difference between the expected bits per pixel (bpp) of compressed ($\text{bpp}_{\text{enc}}$) and uncompressed ($\text{bpp}_{\text{raw}}$) sensings. The raw bpp value refers to the bit depth, i.e., the sensor’s radiometric resolution and the number of bands. For example, the radiometric resolution of Sentinel-2 is 12 bits [49], so for three bands $\text{bpp}_{\text{raw}}=3\cdot 2^{12}$.
The advantage of TCR/s is twofold. First, decoupling it from system-specific parameters, such as sensor resolution or orbital period, permits drawing generalizable insights regarding the relative trade-off between codec overhead and bitrate reduction. Second, operators can still assess the feasibility of the pipeline on target hardware and the expected downlinkable data volume by running the profiler with configurations that reflect deployment conditions.
@@ 4.3 Concurrent Task Execution
So far, this section has solely discussed the computational cost of a codec’s parametric transforms without considering pre- and post-processing. In particular, after applying the encoder transforms, it is still necessary to entropy code the output to compress the latent. Since FOOL’s entropy model is input adaptive, it requires a range coder. Although more recent range coders are efficient, they incur non-negligible runtime overhead. Therefore, given the unforgiving conditions of OEC, we argue that the entropy coder cannot be neglected in the design process and evaluation of a neural codec. FOOL virtually offsets the entire runtime overhead with simple concurrent task execution. The idea is to exploit the minimal interference of processes that draw from different resource types.
Specifically, FOOL offsets virtually the entire runtime overhead with simple concurrent task execution. The idea is to exploit the minimal interference of processes drawn from different resource types. For three sequentially incoming samples $x_{i-1},x_{i},x_{i+1}$, FOOL executes CPU-bound pre-processing of $x_{i+1}$, accelerator-bound inference of $x_{i}$, and CPU-bound post-processing of $x_{i-1}$. In this work, pre-processing corresponds to tiling the samples, and post-processing to entropy coding with rANS [50, 51]. Concurrently to inference ${x_{i-1}}$ on the accelerator, a process starts tiling $x_{i}$. After inference on ${x_{i-1}}$, $\boldsymbol{\hat{\mathrm{y}}},\boldsymbol{\hat{\mathrm{z}}},\boldsymbol{\hat{\sigma}},\boldsymbol{\hat{\mu}}$ (Section 5.7) are persisted on the file system. Then, a separate process loads the data and losslessly compresses $\boldsymbol{\hat{\mathrm{y}}},\boldsymbol{\hat{\mathrm{z}}}$ with an entropy coder. We expect minimal interference between the processes, resulting in virtually no PP/s decrease.
@@ 5 The FOOL’s Compression Method
We design the compression method based on three criteria. First, it should synergize with the profiling strategy (Section 4.2). Second, it should embed context for feature compression without favoring a particular downstream task. Third, it should prioritize machine perception but allow recovering human interpretable images without increasing the bitrate (e.g., with an additional distortion term in the loss function).
@@ 5.1 Model Building Blocks
Figure 6 illustrates the architecture’s building blocks and how it organizes the primary networks for transform coding (Section 3.2.1). The primary networks have four stages that control the depth and width. Each stage has at most one lightweight attention block and at least one residual block. A residual block optionally up or downsamples the spatial dimensions. A stage’s width and depth parameters configure the number of channels and residual blocks within a stage.
@@ 5.2 Capturing Inter-Tile Dependencies
The input to the compression model are tiles that were partitioned to maximize processing throughput (Section 4.2), i.e., we consider an input $\boldsymbol{\mathrm{x}}$ as a list with $T$ separate image tensors $x_{t}\in\mathbb{R}^{C\times H\times W}$.
To decrease transfer costs further, FOOL leverages the prior knowledge from partitioning (i.e., tiles corresponding to the same image) in two ways. The first is via weight-sharing with 2D Residual Blocks by reshaping the tensor to $T\cdot B\times C\times H\times W$. This way, we include further inductive bias during training by forward passing $T$ similar tensors before each backpropagation. The second is with an inter-tile attention mechanism. Since self-attention from transformer architectures is prohibitively expensive for our purposes (even in the latent space as proposed in [52]). Therefore, we modify and extend the lightweight convolutional attention layer from [53]. The layer stacks residual units to increase the receptive field that primarily emphasizes local interactions. FOOL partially replaces the need for global operations by assuming tiles to have “pseudo-temporal” dependencies. Intuitively, partitioning single captures that span large geographic areas may be similar to moving a video feed with large strides. In particular, tiles within the same regions or biomes have global dependencies.
In the 3D version of the attention block from Figure 6, a residual unit consists of two $1\times 1\times 1$ convolution and a $D\times 3\times 3$ convolution in between. The kernel size for the temporal dimension of attention layers (Section 5.1) is set as $D=3$ for $T<5$ and $D=5$ for $T\geq 5$.
The advantage of 3D layers over concatenating channels and applying 2D convolutional operations (e.g., with channel attention [54]) is that it considerably reduces width. For example, given a $3\times 3$ 2D convolution with $T\times C$ in and out channels. Then for $T=5$ image tensors, with dimensions $C=64,H=128,W=128$, would require $5\cdot 64\cdot(3\cdot 3\cdot 5\cdot 64+1)=921,920$ parameters. Conversely, for a $5\times 3\times 3$ 3D convolution with the same in and out channels, it would result in $64\cdot(5\cdot 3\cdot 3\cdot 64+1)=184,384$. Additionally, we reduce the number of multiply-and-accumulates from approximately 15 million to 9 million. This is noteworthy, since besides lowering memory requirements, it allows FOOL to scale model capacity with less friction against processing throughput (Section 4.2).

### v2

@@ 3.3.1 The Effectiveness of Shallow Features
Figure 3 visualizes the trade-off. We extract the features from a ResNet network with weights trained on the ImageNet [38] classification task and recover the original image by training separate reconstruction models for each marked location. Section 5.6 elaborates on the reconstruction. Notice that models at shallow layers can recover the input with high similarity, but the recovery progressively worsens as the path distance increases. Now assume a discriminative model (e.g., a ResNet for image classification) $\mathcal{M}=(\mathcal{H},\mathcal{T})$ with separate shallow $\mathcal{H}$ from deeper layers $\mathcal{T}$ as disjoint subsets, such that $\mathcal{H}(X)=H$ (i.e, mapping input to shallow features) and $\mathcal{M}(X)=\mathcal{T}(\mathcal{H}(X))$. Further, given a codec $\mathrm{c}=(\mathrm{enc},\mathrm{dec})$ where $\mathrm{dec}(\mathrm{enc}(X))=\tilde{H}$ is an approximation of $\mathcal{H}(X)$. Then, $\tilde{H}$ is a sufficient approximation of $H$ if $\mathcal{T}(\tilde{H})$ results in lossless prediction, i.e., no drop in prediction performance relative to $\mathcal{T}(H)$. In other words, a sufficient representation in the shallow latent space results in high similarity in the deep latent space between $\mathcal{T}(H)$ and $\mathcal{T}(\tilde{H})$. Consider that similarity in the deep latent space coincides with high similarity for human perception in the input space [18]. Therefore, the encoder output $\texttt{enc}(X)$ should retain sufficient information to reconstruct $X$ with quality comparable to $\mathcal{H}(X)$, as exemplified in Figure 3. Finally, since $\texttt{enc}(X)$ sufficiently approximates $H$, it should be possible to sufficiently approximate the shallow representation of any model $\mathcal{M}^{\prime}=(\mathcal{H}^{\prime},\mathcal{T}^{\prime})$ if $I(\mathcal{H}(X),X)\approx I(\mathcal{H}^{\prime}(X),X)$.
3.3.2 Rate Reductions by Task Specificity
The idea of task-oriented communication is that messages for model prediction may require less information than human domain experts, i.e., that it should be possible to reduce bitrate by not (exclusively) using input reconstruction as the distortion measure. We argue that this assumption contradicts empirical evidence demonstrating models outperforming human experts in various image-related tasks, i.e., machines can detect signals and patterns that humans physiologically or intellectually cannot. Rather, the opposite should hold, i.e., when compressing for quality using domain experts as judges, we should see more potential for rate savings, not less. The claim is consistent with the results in Figure 1 where codecs with high reconstruction quality result in images that are deceptively similar to the input (details in Section 6.3). Conclusively, rate reductions are from task specificity of the distortion measure, irrespective of the input interface, whether it is a particular layer of a DNN architecture, human receptors, or textual encoding. Note that this hols, even if when limiting measures to discriminative task objectives without any image reconstruction. Besides visualizing Equation 8, the input image illustrates a practical example. The frog subset of ImageNet distinguishes between Tree Frogs, Bullfrogs, and Tailed Frogs. Since these frog species have distinct figures and dominant colors, the more delicate characteristics of a tree frog are redundant for ImageNet classification. The network gradually discards information regarding the fine-grained blue-yellow colored patterns, permitting only the recovery of general shape and environment from the deep features. The deeper the features, the less structure and detail are present, which may be redundant for the task. Now, suppose training a codec where the encoder retains the minimal information necessary to reconstruct the output of the deepest layers for a classification task (e.g., similar to Vista [24]). Then, we can reduce the transfer cost to as low as $\log_{2}{({\#\text{labels}})}$ without predictive loss. However, we may lack the information for other tasks, i.e., there is a trade-off between generalization and the lower bound on the bitrate. In contrast, targeting shallow features for compression may strike a balance between aiming to retain information for all possible downstream tasks and only emphasizing the salient regions for the tasks associated with a foundational model. Arguably, the limitation is negligible, as maintainers will train foundational models with useful tasks in mind.
@@ 4.1 Compression and Prediction Request Flow
Figure 4 illustrates a high-level view for serving requests.
@@
For samples processed by FOOL, there is a single encoder. The output $\boldsymbol{\hat{{\mathrm{y}}}}$ is forwarded to the detection pipeline, skipping the shallow layers. A decoder transforms $\boldsymbol{\hat{{\mathrm{y}}}}$ into a target representation. A predictor outputs bounding boxes for a specified task. An image reconstruction model optionally restores the latent to a human-interpretable image to overlay the bounding boxes. Samples downlinked with bent pipe or some image codec are forwarded to the shallow layers, skipping the FOOL decoder and reconstruction model. This section focuses on the pipeline, before Section 5 introduces the compression method.
@@ 4.2 Profiling Compression Pipelines for OEC
A common challenge for operators is to determine whether reported performance regarding resource usage or throughput from the latest advancements generalizes to their target hardware. This problem stems not from a lack of rigor by authors but from the sheer heterogeneity of the AI accelerator landscape [48]. Graph compilers and other vendor-specific optimizations (e.g., TensorRT [^2] https://developer.nvidia.com/tensorrt, Apache TVM [^3] https://tvm.apache.org) further complicate evaluation, with varying methods for operator fusion, graph rewriting, etc. Consequently, FOOL includes a simple profiling and evaluation strategy that operators may run before deployment. Notably, in contrast to existing work that partitions images to match the input size of a particular application, the profiler regards the importance of spatial dimensions for resource efficiency. The purpose of the profiler is to determine a configuration that maximizes throughput. While throughput evaluation is straightforward, how to measure it (e.g., images/second) is not necessarily obvious, particularly for (neural) compression pipelines.
First, terrestrial and LEO remote sensing with constrained sensor networks demand resource-conscious methods, but in LEO, downlinks are only available within segments. Due to memory and storage constraints, devices must process samples according to a sensor rate, i.e., a prolonged interval between incoming samples. Hence, the objective in LEO is to maximize the number of pixels the accelerator can process before reaching a downlink segment, given a time constraint for a single sample (i.e., “frame deadline” [27]). For example, assume a cheaper and a costlier compression model where both models meet the frame deadline. Applying the latter results in half the bitrate but thrice the inference time. Using the former is beneficial in most network conditions for real-time terrestrial applications since it results in a lower end-to-end request latency. In contrast, applying the latter in LEO may be advantageous, as finishing earlier results in the needless idle time of resources. Second, satellite imagery has substantially higher resolution than captures from most terrestrial sensor networks. A standard method to improve throughput for high-dimensional images is parallel processing with tile partitioning. The distinction is that there is more control over the spatial dimensions and the batch size. Nonetheless, a caveat is the friction between a model’s size and the input size. Increasing the width (e.g., the number of feature maps output by a convolutional layer) of a neural codec’s parametric transforms may result in better compression performance but lower processing throughput. In summary, we require a measure that includes (i) the tile spatial dimensions, (ii) batch size, and (iii) the capacity-compression performance trade-off.
@@
Notice how evaluating img/s always favors smaller spatial dimensions and disregards batch size and model width. In contrast, PP/s reveals that the optimal spatial dimension is around $500\times 500$ but will naturally favor smaller models, as it does not consider that wider models may reduce transfer costs. To alleviate the limitations of PP/s, we measure Transfer Cost Reduction per Second (TCR/s) as:
@@
The measure now includes the compression performance as the difference between the expected bits per pixel (bpp) of compressed ($\text{bpp}_{\text{enc}}$) and uncompressed ($\text{bpp}_{\text{raw}}$) sensings. The raw bpp value refers to the bit depth, i.e., the sensor’s radiometric resolution and the number of bands. For example, the radiometric resolution of Sentinel-2 is 12 bits [28], so for three bands $\text{bpp}_{\text{raw}}=3\cdot 2^{12}$. The advantage of TCR/s is twofold. First, decoupling it from system-specific parameters, such as sensor resolution or orbital period, permits drawing generalizable insights regarding the relative trade-off between codec overhead and bitrate reduction. Second, operators can still assess the feasibility of the pipeline on target hardware and the expected downlinkable data volume by running the profiler with configurations that reflect deployment conditions.
@@ 4.3 Concurrent Task Execution
So far, this section has solely discussed the computational cost of a codec’s parametric transforms without considering pre- and post-processing. In particular, after applying the encoder transforms, it is still necessary to entropy code the output to compress the latent. Since FOOL’s entropy model is input adaptive, it requires a range coder. Although more recent range coders are efficient, they incur non-negligible runtime overhead. Therefore, given the unforgiving conditions of OEC, we argue that the entropy coder cannot be neglected in the design process and evaluation of a neural codec. FOOL virtually offsets the entire runtime overhead with simple concurrent task execution. The idea is to exploit the minimal interference of processes that draw from different resource types. For three sequentially incoming samples $x_{i-1},x_{i},x_{i+1}$, FOOL executes CPU-bound pre-processing of $x_{i+1}$, accelerator-bound inference of $x_{i}$, and CPU-bound post-processing of $x_{i-1}$. In this work, pre-processing corresponds to tiling the samples, and post-processing to entropy coding with rANS [49, 50]. Concurrently to inference ${x_{i-1}}$ on the accelerator, a process starts tiling $x_{i}$. After inference on ${x_{i-1}}$, $\boldsymbol{\hat{\mathrm{y}}},\boldsymbol{\hat{\mathrm{z}}},\boldsymbol{\hat{\sigma}},\boldsymbol{\hat{\mu}}$ (Section 5.7) are persisted on the file system. Then, a separate process loads the data and losslessly compresses $\boldsymbol{\hat{\mathrm{y}}},\boldsymbol{\hat{\mathrm{z}}}$ with an entropy coder. We expect minimal interference between the processes, resulting in virtually no PP/s decrease.
@@ 5 The FOOL’s Compression Method
We design the compression method based on three criteria. First, it should synergize with the profiling strategy (Section 4.2). Second, it should embed context for feature compression without favoring a particular downstream task. Third, it should prioritize the integrity of downstream tasks but allow recovering human interpretable images without increasing the bitrate.
@@ 5.1 Model Building Blocks
Figure 6 illustrates the architecture’s building blocks and how it organizes the primary networks for transform coding (Section 3.2). The primary networks have four stages that control the depth and width. Each stage has at most one lightweight attention block and at least one residual block. A residual block optionally up or downsamples the spatial dimensions. A stage’s width and depth parameters configure the number of channels and residual blocks within a stage.
@@ 5.2 Capturing Inter-Tile Dependencies
The input to the compression model are tiles that were partitioned to maximize processing throughput (Section 4.2), i.e., we consider an input $\boldsymbol{\mathrm{x}}$ as a list with $T$ separate image tensors $x_{t}\in\mathbb{R}^{C\times H\times W}$. To decrease transfer costs further, FOOL leverages the prior knowledge from partitioning (i.e., tiles corresponding to the same image) in two ways. The first is via weight-sharing with 2D Residual Blocks by reshaping the tensor to $T\cdot B\times C\times H\times W$. This way, we include further inductive bias during training by forward passing $T$ similar tensors before each backpropagation. The second is with an inter-tile attention mechanism. Since self-attention from transformer architectures is prohibitively expensive for our purposes, even when applying it on downsampled representations as proposed in [51]. Therefore, we modify and extend the lightweight convolutional attention layer from [52]. The layer stacks residual units to increase the receptive field that primarily emphasizes local interactions. FOOL partially replaces the need for global operations by assuming tiles to have “pseudo-temporal” dependencies. Intuitively, partitioning single captures that span large geographic areas may be similar to moving a video feed with large strides. In particular, tiles within the same regions or biomes have global dependencies. In the 3D version of the attention block from Figure 6, a residual unit consists of two $1\times 1\times 1$ convolution and a $D\times 3\times 3$ convolution in between. The kernel size for the temporal dimension of attention layers (Section 5.1) is set as $D=3$ for $T<5$ and $D=5$ for $T\geq 5$.
The advantage of 3D layers over concatenating channels and applying 2D convolutional operations (e.g., with channel attention [53]) is that it considerably reduces width. For example, given a $3\times 3$ 2D convolution with $T\times C$ in and out channels. Then for $T=5$ image tensors, with dimensions $C=64,H=128,W=128$, would require $5\cdot 64\cdot(3\cdot 3\cdot 5\cdot 64+1)=921,920$ parameters. Conversely, for a $5\times 3\times 3$ 3D convolution with the same in and out channels, it would result in $64\cdot(5\cdot 3\cdot 3\cdot 64+1)=184,384$. Additionally, we reduce the number of multiply-and-accumulates from approximately 15 million to 9 million. Besides lowering memory requirements, this allows FOOL to scale model capacity with less friction against processing throughput.